# INR-TH-2025-001

### v1

@@ Abstract
However, $\theta$ may also be limited from below if the only source of sterile neutrino DM production is the sterile-active mixing  [^1] There are many models where other interactions are producing DM sterile neutrinos [11, 12, 13, 14, 15]. It is this case which will be studied below.
@@
The lepton asymmetry may influence the later cosmological stages, in particular, primordial nucleosynthesis and recombination. Observations of primordial helium abundance favor standard cosmology and severely constrain [16] the asymmetry in electron neutrinos at the moment of neutron freeze-out, which happens at the plasma temperature of about $T_{n}\approx 0.7$ MeV, some time after neutrino decoupling. As we show below, the lower bounds on mixing $\theta$, inferred from this constraint, can be significantly reduced in the case of a specific initial lepton asymmetry pattern, with large individual flavor asymmetries but a total lepton asymmetry close to zero. Then active neutrino oscillations, which start at temperature $T\lesssim 20$ MeV, very effectively cancel all flavor asymmetries in accordance with calculations of Ref. [17]. We find that the cancellation is more efficient the in case of inverse ordering of neutrino masses and CP-phase does not play a noticeable role. This allows the larger flavor lepton asymmetries at sterile neutrino production, hence lower mixing angles, consistent with observed helium abundance and anisotropy of the cosmic microwave background.
@@
Note that the production rate of sterile neutrinos grows proportionally to the weak interaction rate, see r.h.s. of eq. (1), since interaction in plasma collapses the neutrino wave function, which results in the appearance of sterile neutrino with probability (2). The weak rate is higher at higher temperature, but the transition probability (2) is highly suppressed both by damping (the second term in denominator) and by thermal potential. Hence, generically the oscillation probability in plasma is lower than in vacuum, but for the lepton asymmetry potential which may cancel the mass term in the brackets and thus amplifies the transition similarly to the Mikheev–Smirnov–Wolfenstein effect in the Sun. The lepton asymmetry potential also decreases with temperature, $V^{L}_{\alpha}\propto T^{3}$, but slower than the thermal potential $V^{T}_{\alpha}\propto T^{5}$. Therefore, the oscillation probability is maximized when the quantity in brackets is close to zero,
@@
Figure 1: Spectra of sterile neutrinos produced via mixing with electron neutrinos, $y=p/T$, at various plasma temperatures $T$ and various initial lepton asymmetries. The initial flavor neutrino asymmetries are equal to $(1,\,-0.5\,-0.5)\cdot N\cdot 10^{-3}$ for electron, muon and tau. The initial values of the scale factor $N$ used in calculations for the three plots are 1, 3,91391,\,3\,,91 , 3 , 9 from left to right.
@@
Second, the resonance momentum becomes colder with higher asymmetry $\Delta_{\alpha}$. It is confirmed in Fig. 1 with numerical calculations for three different initial asymmetries. At the same temperature the spectra are colder for higher asymmetries (from left to right on the plot) in full accordance with (6). Hence, exploiting higher asymmetry also helps to avoid the problems with cosmic structure formation and reopens a window of light sterile neutrinos. Moreover, with higher asymmetry the production lasts longer, reaching the epoch of active neutrino oscillations, $T\gtrsim 20$ MeV which very efficiently cancel asymmetry between the neutrino flavors (see below) hence finishing the sterile neutrino production and consequently yielding generically colder spectrum.
Alternatively, one may reduce the asymmetry and stop the production by involving some new ingredients. There is an issue here. Though, the presented in Fig. 1 spectra clearly show what can be achieved with instant disappearance of the lepton asymmetry, in realistic mechanisms the operation takes some time, and so the spectra are smoother than one could envisage from the plots. Actually, by reducing the asymmetry one shifts the most effective production to earlier stages but makes the production longer, and so the average momentum remains almost intact. We observed this behavior in a realistic situation with one more, but much heavier sterile neutrino component in the plasma. We assume, that these neutrinos are non-relativistic and their decay rate is about the Hubble parameter in the interesting epoch, $\Gamma_{d}=H(T_{d})$. The decay products are asymmetric in neutrino flavors, and tends to reduce the neutrino asymmetries. This reduction changes the final spectra of sterile neutrino very little while we vary the temperature in the interval $26080<T_{d}<$ 260 MeV.
@@
Figure 2: Spectra of sterile neutrinos produced via mixing with electron neutrinos, $y=p/T$, at various plasma temperatures $T$ and initial lepton asymmetries $(1,\,-0.5\,-0.5)\cdot 10^{-3}$ for electron, muon and tau. The asymmetries are reduced because of flavor violating heavy neutral fermions decays of heavy neutral fermions in plasma with decay rates $\Gamma_{d}=H(T_{d})\times(T_{d}/T)^{10}$ and reference temperatures $T_{d}$ (from left to right): 80 MeV, 100 MeV, 130 MeV.
One can see that the maxima of the final spectra shift towards the cold direction with earlier destruction of the lepton asymmetry (from possible decays of heavy fermions). There is no need to make the asymmetry equal to zero, just to reduce it significantly, e.g. for the right plot in Fig. 2 the final neutrino asymmetries are (1.51,1.75, $(1.51,\,-1.75,\,2.03)\cdot 10^{-4}$.
3. The larger lepton asymmetry yields the strongest amplification of the resonance production. One can check that at large asymmetry the amount of produced particles scales linearly with asymmetry. Thus, the larger the asymmetry, the lower the mixing angle $\theta$ is needed to provide the same amount of dark matter. However, when oscillations into sterile neutrino terminate, the net lepton asymmetry remains in plasma. In the literature, typically, it is supposed that the asymmetry resides mostly in one neutrino flavor. Then, the active neutrino oscillations in plasma, which begin at temperature of about 20 MeV, redistribute it over all three species. Then the electron neutrino asymmetry approaches one third of the initial asymmetry and must obey the upper limit inferred from observations of the helium abundance [16]. It implies a lower limit on the mixing angle $\theta$ consistent with the hypothesis that sterile neutrinos, generated in this way, fully explain the dark matter component.
@@
Potentially, this observation allows one to use any large asymmetry at the beginning of sterile neutrino production as far as the sum of final lepton flavor asymmetries is close to zero. In practice, the lepton asymmetry redistribution is quite limited in time: it starts with onset of oscillations at temperature of about 20 MeV and stops with neutrino decoupling from primordial plasma, when its temperature drops to about 3 MeV.
@@
we illustrate this chain of reasoning by numerical calculations of sterile neutrino production performed with dmpheno code [21] for $m_{\nu_{s}}=7.11$ keV, mixing with electron neutrino, $\sin^{2}2\theta=7\cdot 10^{-14}$, and two different sets of initial neutrino flavor asymmetries, $(\Delta_{\nu_{e}},\Delta_{\nu_{\mu}},\Delta_{\nu_{\tau}})$, namely $(2,-1,-1)\cdot 10^{-4}$ (left panel) and $(50,-25,-25)\cdot 10^{-4}$ (right panel) and subsequent evolution of the neutrino asymmetries at the stage of active neutrino oscillations, performed with FortEPiaNO code [22] and present estimates of the neutrino oscillation parameters [24]. The values of interest are the ratios of the neutrino chemical potentials to the temperature $\xi_{\nu_{\alpha}}\equiv\mu_{\nu_{\alpha}}/T$. They are related to the neutrino asymmetries $\Delta_{\nu_{\alpha}}$ used above as follows,
@@
We find that the interesting electron neutrino asymmetry starts to reduce quite late indeed, but the efficiency of the asymmetry reduction differs in the two cases of mass hierarchy due to different combination of the mixing angles determining the numerical factor $F$ in the reduction exponent of (7). In case of normal hierarchy (solid lines in Fig. 3) the reduction factor is only about ten. Since numerically $\xi_{\nu_{e}}\approx 33\Delta_{\nu_{e}}$, with the present limit from BBN $|\xi_{\nu_{e}}|<0.01$ [16] one may use the initial electron neutrino asymmetry as large as $\Delta_{\nu_{e}}\simeq 10^{-3}$. In case of inverse hierarchy (dotted lines in Fig. 3) the reduction factor is much larger, and one may adopt much higher initial neutrino asymmetry.
@@
5. To summarize, in this letter, we show that the presently allowed model parameter space of the sterile neutrino dark matter may be significantly extended with specific tuned neutrino flavor asymmetries. If large in magnitude, but the sum is close to zero, it can be efficiently cancel in the subsequent oscillations of the active neutrino. The inverse ordering of neutrino masses shows more efficient cancellation. Hence, at the moment of neutron decoupling from plasma the electron asymmetry may be greatly reduced, even to zero. It mitigates the limits from primordial nucleosynthesis, allows for higher initial lepton asymmetry and hence lower sterile-neutrino mixing angles. Likewise, the with higher initial lepton asymmetry the resonant mechanism produces colder sterile neutrinos first. We show, that if the lepton asymmetry disappears from the plasma, the resulting spectrum is indeed colder. At the same time, higher asymmetry postpones the sterile neutrino production, and with sufficiently high asymmetry the active neutrino oscillations may do the job, cancel the asymmetry and make the sterile neutrino spectrum cold.
Note in passing that some electron asymmetry at the level of few percent may be welcome given some anomalous abundances of primordial elements recently observed, see e.g.[16].
It is worth to mention, that while there are other numerical codes to calculate the neutrino oscillations in the primordial plasma, we believe that the qualitative results do not depend on our choice. The qualitative picture with various ways to reopen the areas of the model parameter space is correct.
Without using a particular mechanism to produce the original lepton asymmetry, one is quite free to tune the parameters in the sterile neutrino sector. We illustrate our statements with numerical calculations performed for particular values of sterile neutrino mass and mixing angle. The parameters of the active neutrino sector are fixed by neutrino oscillation experiments, except the mass hierarchy and the value of CP-phase. As we observed above, the normal hierarchy is capable of reducing the net neutrino asymmetry more efficiently. We fix the value of CP-phase as suggested in [24], but the present accuracy of it’s determination is poor. Thus we run the FortEPiaNO code with the suggested value of CP-phase and with zero CP-phase for the set of neutrino asymmetries and do not find any significant change in the dynamics of asymmetry, see Fig.,5.

### v2

@@ Abstract
However, $\theta$ may also be limited from below if the only source of sterile neutrino DM production is the sterile-active mixing  [^1] There are many models where other interactions are producing DM sterile neutrinos [11, 12, 13, 14, 15].. It is this case which we study below.
@@
The lepton asymmetry may influence the later cosmological stages, in particular, primordial nucleosynthesis and recombination. Observations of primordial helium abundance favor standard cosmology and severely constrain [16, 17] the asymmetry in electron neutrinos at the moment of neutron freeze-out, which happens at the plasma temperature of about $T_{n}\approx 0.7$ MeV, some time after neutrino decoupling. As we show below, the lower bounds on mixing $\theta$, inferred from this constraint, can be significantly reduced in the case of a specific initial lepton asymmetry pattern, with large individual flavor asymmetries but a total lepton asymmetry close to zero. Then active neutrino oscillations, which start at temperature $T\lesssim 20$ MeV, very effectively cancel all flavor asymmetries in accordance with calculations of Ref. [18]. We find that the cancellation is more efficient the in case of inverse ordering of neutrino masses and CP-phase does not play a noticeable role. This allows the larger flavor lepton asymmetries at sterile neutrino production, hence lower mixing angles, consistent with observed helium abundance and anisotropy of the cosmic microwave background.
@@
Note that the production rate of sterile neutrinos grows proportionally to the weak interaction rate, see r.h.s. of eq. (1), since interaction in plasma collapses the neutrino wave function, which results in the appearance of sterile neutrino with probability (2). The weak rate is higher at higher temperature, but the transition probability (2) is highly suppressed both by damping (the second term in the denominator) and by thermal potential. Hence, generically the oscillation probability in plasma is lower than in vacuum, but for the lepton asymmetry potential which may cancel the mass term in the brackets and thus amplifies the transition similarly to the Mikheev–Smirnov–Wolfenstein effect in the Sun. The lepton asymmetry potential also decreases with temperature, $V^{L}_{\alpha}\propto T^{3}$, but slower than the thermal potential $V^{T}_{\alpha}\propto T^{5}$. Therefore, the oscillation probability is maximized when the quantity in brackets is close to zero,
@@
Figure 1: Spectra of sterile neutrinos produced via mixing with electron neutrinos, $y=p/T$, at various plasma temperatures $T$ and various initial lepton asymmetries. The initial flavor neutrino asymmetries are equal to $(1,\,-0.5\,-0.5)\cdot N\cdot 10^{-3}$ for electron, muon and tau. The initial values of the scale factor $N$ used in calculations for the three plots are 1, 3,91391,\,3\,,91 , 3 , 9 from left to right. The dips in the spectrum are caused by the limited precision of the numerical calculations, which could be improved but at the cost of significantly increased computational time.
@@
Second, the resonance momentum becomes colder with higher asymmetry $\Delta_{\alpha}$. It is confirmed in Fig. 1 with numerical calculations for three different initial asymmetries. At the same temperature, the spectra are colder for higher asymmetries (from left to right on the plot) in full accordance with (6). Hence, exploiting higher asymmetry also helps to avoid the problems with cosmic structure formation and reopens a window of light sterile neutrinos. Moreover, with higher asymmetry the production starts later, reaching the epoch of active neutrino oscillations, $T\gtrsim 20$ MeV, which very efficiently cancel asymmetry between the neutrino flavors (see below) hence finishing the sterile neutrino production and consequently yielding generically colder spectrum.
Alternatively, one may reduce the asymmetry and stop the production by involving some new ingredients. There is an issue here. Though, the presented in Fig. 1 spectra clearly show what can be achieved with instant disappearance of the lepton asymmetry, in realistic mechanisms the operation takes some time, and so the spectra are smoother than one could envisage from the plots. Actually, by reducing the asymmetry one shifts the most effective production to earlier stages but makes the production longer, and so the average momentum remains almost intact. We observed this behavior in a realistic situation with one more, but much heavier sterile neutrino component in the plasma. We assume, that these neutrinos are non-relativistic and their decay rate is about the Hubble parameter in the interesting epoch, $\Gamma_{d}=H(T_{d})$. The decay products are asymmetric in neutrino flavors and tend to reduce the neutrino asymmetries. This reduction changes the final spectra of the sterile neutrino very little, while we vary the temperature in the interval $26080<T_{d}<$ 260 MeV.
@@
Figure 2: Normalized to unity spectra of sterile neutrinos produced via mixing with electron neutrinos, $y=p/T$, at various plasma temperatures $T$ and initial lepton asymmetries $(1,\,-0.5\,-0.5)\cdot 10^{-3}$ for electron, muon and tau. The asymmetries are reduced because of flavor violating heavy neutral fermions decays of heavy neutral fermions in plasma with decay rates $\Gamma_{d}=H(T_{d})\times(T_{d}/T)^{10}$ and reference temperatures $T_{d}$ (from left to right): 80 MeV, 100 MeV, 120 MeV.
One can see that the maxima of the final spectra shift towards the cold direction with earlier destruction of the lepton asymmetry (from possible decays of heavy fermions). There is no need to make the asymmetry equal to zero, just to reduce it significantly, e.g. for the right plot in Fig. 2 the final neutrino asymmetries are (1.93,0.04, $(-1.93,\,-0.04,\,3.75)\cdot 10^{-4}$. Reducing the asymmetry in this manner shifts the resonant production toward higher values of y, which results in a warmer spectrum. This effect competes with the cooling effect of the decays, which consequently limits the reduction in the mean momentum $p=yT$).
3. The larger lepton asymmetry yields the strongest amplification of the resonance production. One can check that at large asymmetry the amount of produced particles scales linearly with asymmetry. Thus, the larger the asymmetry, the lower the mixing angle $\theta$ is needed to provide the same amount of dark matter. However, when oscillations into sterile neutrino terminate, the net lepton asymmetry remains in plasma. In the literature, typically, it is supposed that the asymmetry resides mostly in one neutrino flavor. Then, the active neutrino oscillations in plasma, which begin at temperature of about 20 MeV, redistribute it over all three species. Then the electron neutrino asymmetry approaches one third of the initial asymmetry and must obey the upper limit inferred from observations of the helium abundance [16, 17]. It implies a lower limit on the mixing angle $\theta$ consistent with the hypothesis that sterile neutrinos, generated in this way, fully explain the dark matter component.
@@
Potentially, this observation allows one to use any large asymmetry at the beginning of sterile neutrino production as far as the sum of final lepton flavor asymmetries is close to zero. In practice, the lepton asymmetry redistribution is quite limited in time. It starts with the onset of oscillations at a temperature of about 20 MeV and stops with neutrino decoupling from the primordial plasma, when its temperature drops to about 3 MeV.
@@
we illustrate this chain of reasoning by numerical calculations of sterile neutrino production performed with dmpheno code [22] for $m_{\nu_{s}}=7.11$ keV, mixing with electron neutrino, $\sin^{2}2\theta=7\cdot 10^{-14}$, and two different sets of initial neutrino flavor asymmetries, $(\Delta_{\nu_{e}},\Delta_{\nu_{\mu}},\Delta_{\nu_{\tau}})$, namely, $(2,-1,-1)\cdot 10^{-4}$ (left panel) and $(50,-25,-25)\cdot 10^{-4}$ (right panel) and subsequent evolution of the neutrino asymmetries at the stage of active neutrino oscillations, performed with FortEPiaNO code [23] and present estimates of the neutrino oscillation parameters [16]. The values of interest are the ratios of the neutrino chemical potentials to the temperature $\xi_{\nu_{\alpha}}\equiv\mu_{\nu_{\alpha}}/T$. They are related to the neutrino asymmetries $\Delta_{\nu_{\alpha}}$ used above as follows,
@@
We find that the interesting electron neutrino asymmetry starts to reduce quite late indeed, but the efficiency of the asymmetry reduction differs in the two cases of mass hierarchy due to different combination of the mixing angles determining the numerical factor $F$ in the reduction exponent of (7). In case of normal hierarchy (solid lines in Fig. 3) the reduction factor is only about ten. Since numerically $\xi_{\nu_{e}}\approx 33\Delta_{\nu_{e}}$, with the present limit from BBN $|\xi_{\nu_{e}}|\lesssim 0.01$ [16, 17] one may use the initial electron neutrino asymmetry as large as $\Delta_{\nu_{e}}\simeq 10^{-3}$. In case of inverse hierarchy (dotted lines in Fig. 3) the reduction factor is much larger, and one may adopt much higher initial neutrino asymmetry.
@@
5. To summarize, in this letter we show that the presently allowed model parameter space of the sterile neutrino dark matter may be significantly extended with specific tuned neutrino flavor asymmetries. If large in magnitude, but the sum is close to zero, it can be efficiently cancel in the subsequent oscillations of the active neutrino. The inverse ordering of neutrino masses shows more efficient cancellation. Hence, at the moment of neutron freeze-out the electron asymmetry may be greatly reduced, even to zero. It mitigates the limits from primordial nucleosynthesis, allows for higher initial lepton asymmetry and hence lower sterile-neutrino mixing angles. Likewise, with a higher initial lepton asymmetry the resonant mechanism produces colder sterile neutrinos. We show, that if the lepton asymmetry disappears from the plasma, the resulting spectrum is indeed colder. At the same time, higher asymmetry postpones the sterile neutrino production, and with sufficiently high asymmetry the active neutrino oscillations may do the job, cancel the asymmetry and make the sterile neutrino spectrum cold.
Note in passing that some electron asymmetry at the level of few percent may be welcome given some anomalous abundances of primordial elements recently observed, see e.g.[25, 17].
It is worth mentioning that while there are other numerical codes to calculate the neutrino oscillations in the primordial plasma, we believe that the qualitative results do not depend on our choice. The qualitative picture with various ways of reopening the areas of the model parameter space is correct.
Without using a particular mechanism to produce the original lepton asymmetry, one is quite free to tune the parameters in the sterile neutrino sector. We illustrate our statements with numerical calculations performed for particular values of sterile neutrino mass and mixing angle. The parameters of the active neutrino sector are fixed by neutrino oscillation experiments, except the mass hierarchy and the value of CP-phase. As we observed above, the inverted hierarchy is capable of reducing the net neutrino asymmetry more efficiently. We fix the value of CP-phase as suggested in [16], but the present accuracy of it’s determination is poor. Thus we run the FortEPiaNO code with the suggested value of CP-phase and with zero CP-phase for the set of neutrino asymmetries and do not find any significant change in the dynamics of asymmetry, see Fig.,5.